The most usually-studied capacitors is made up of two parallel conducting plates. These are separated by 1.77 mm. What is the capacitance in air if the plate in 17 cm2? What is the charge is a potential difference of 120 V is applied? 


1
Expert's answer
2022-03-25T09:10:55-0400

The capacitance of the parallel plate capacitor

"C=\frac{\epsilon_0A}{d}=\frac{8.85*10^{-12}*17*10^{-4}}{1.77*10^{-3}}\\
=8.5*10^{-12}\:\rm F=8.5\: pF"

The charge at the capacitor plates

"q=CV=8.5*10^{-12}*120\\
=1.02*10^{-9}\:\rm C=1.02\: nC"